The average number of miligrams (mg) of cholesterol in a cup of a certain brand of ice cream is 660 mg, the standard deviation is 35 mg. Assume the variable is normally distributed.



A. If a cup of ice cream is selected, what is the probability that the cholesterol content will be more than 670 mg?



B. If a sample of 10 cups of ice cream is selected, what is the probability that the mean of the sample will be larger that 670 mg?

1
Expert's answer
2022-05-23T10:53:37-0400

Given "\mu=660\ mg, \sigma=35\ mg, n=10"

"Z=\\frac {X-\\mu}{\\sigma\/\\sqrt{n}}\n\u200b"

A. "P(X>670)=P(Z>\frac{670-660}{35})=P(Z>0.29)=1-P(Z<0.29)=0.3859."

(We found P using z-score table)

The probability that the cholesterol content will be more than 670 mg is 0.3859.

B.


"P(X>670)=1-P(X\leq670)""=1-P(Z\leq\dfrac{670-660}{35/\sqrt{10}})""\approx1-P(Z\leq0.9035)\approx0.1831"


(We found P using z-score table)

The probability that the mean of the sample will be larger than 670 mg is 0.1831.